Find the particular solution of each of the differential equation expressing y explicitly in terms of x.

e) dy/dx - 2xy = 0 y=3 when x = 0

f) xdy/dx + y 2= 0 (x>0) y=1/2 when x= 1


1
Expert's answer
2021-09-20T08:36:47-0400

"1\\
\frac{dy}{dx}-2xy=0\\
\frac{dy}{dx}=2xy\\
\frac{dy}{y}=2xdx\\
\text{Integrate both sides}\\
\ln y=x^2+C\\
y=Ae^{x^2}\\
\text{When } x=0, y=3\\
3=Ae^0\\
A=3\\
\text{The particular solution is }\\
y=3e^{x^2}.\\\\

2.\\
x\frac{dy}{dx}+y^2=0\\
xdy=-y^2dx\\
-\frac{dy}{y^2}=\frac{dx}{x}\\
\text{Integrate both sides}\\
\frac{1}{y}=\ln x+ C\\
y=\frac{1}{\ln x+ C}\\
\text{When } x=1, y= \frac{1}{2}\\
\frac{1}{2}=\frac{1}{\ln x+ C}\\
\frac{1}{2}=\frac{1}{C}\\
C=2\\
\text{The particular solution is }\\
y=\frac{1}{\ln x+2}"